OpenStudy (anonymous):

A positive charge of 0.800μ C is located in a uniform field of 9.50×104 N/C. A negative charge of -0.400μ C is brought near enough to the positive charge that the attractive force between the charges just equals the force on the positive charge due to the field. How close are the two charges?

OpenStudy (jamesj):

The force on the charge \( q \) due to the electric field is \[ F_{field} = Eq \] And the force on the charge \( q \) due to the second charge \( q' = -0.4 \mu C \) is the Coulomb force, \[ F_c = \frac{kqq'}{r^2} \] where \( r \) is the distance you need to calculate. Set the two charges equal and solve for \( r \).

OpenStudy (jamesj):

*two FORCES equal and solve for \( r \)

OpenStudy (anonymous):

do i find the absolute value of q' ? because I cant take the sqrt of a neg #

OpenStudy (jamesj):

The absolute value of the two forces are equal, yes

OpenStudy (anonymous):

I can't get the correct answer for some reason

OpenStudy (jamesj):

If there two are equal, then \[ Eq = kqq'/r^2 \ \implies r = \sqrt{E/k|q'|} \]

OpenStudy (jamesj):

absolute value of the two => ... Just make sure everything is in SI units.

OpenStudy (anonymous):

sqrt of (9.5*10^4)*(9*10^9)*(.4*10^-6) is that right?

OpenStudy (anonymous):

Going by the last formula that James provided, don't you need to divide the value for the electric field by k and q before squaring?

OpenStudy (anonymous):

I'm confused

OpenStudy (jamesj):

First, do you agree with the algebraic derivation above that \[ r = \sqrt{\frac{E}{k|q'|}} \]

OpenStudy (anonymous):

You should be calculating the following quantity\[r= \sqrt( 9.5\times10^{4} )/(9 \times10^{9} \times (0.4*10^{-6})) \]

OpenStudy (anonymous):

I got 5.14 for that answer

OpenStudy (anonymous):

I get that too. Why isn't it correct?

OpenStudy (jamesj):

What's the official answer?

OpenStudy (anonymous):

I am not sure it just tells me when i enter it into lon capa that it is wrong

OpenStudy (jamesj):

Oh, I see. There's an algebraic error ... which one of you should have spotted as well!! ;-) It is that r should be the reciprocal of that expression. So r is \[ r = 0.195 \ m \] Try that

OpenStudy (jamesj):

Assuming everything is positive, if \[ Eq = kqq'/r^2 \] then \[ r = \sqrt{kq'/E} \]

OpenStudy (anonymous):

nice pick up

OpenStudy (anonymous):

that was correct! my brain so so fried right now math is not my strengh. thank you guys for helping me through this